\begin{thebibliography}{9}

%\cite{Caswell:1985ui}
\bibitem{Caswell:1985ui}
W.~E.~Caswell and G.~P.~Lepage,
%``Effective Lagrangians For Bound State Problems 
%In QED, QCD, And Other Field Theories,''
Phys.\ Lett.\ B {\bf 167} (1986) 437;
%%CITATION = PHLTA,B167,437;%%
%\bibitem{Bodwin:1994jh}
G.~T.~Bodwin, E.~Braaten and G.~P.~Lepage,
%``Rigorous QCD analysis of inclusive annihilation 
%and production of heavy quarkonium,''
Phys.\ Rev.\ D {\bf 51} (1995) 1125
[Erratum-ibid.\ D {\bf 55} (1997) 5853]
[hep-ph/9407339].
%%CITATION = HEP-PH 9407339;%%

%\cite{Pineda:1997bj}
\bibitem{Pineda:1997bj}
A.~Pineda and J.~Soto,
%``Effective field theory for ultrasoft momenta in NRQCD and NRQED,''
Nucl.\ Phys.\ Proc.\ Suppl.\  {\bf 64} (1998) 428
[hep-ph/9707481];
%%CITATION = HEP-PH 9707481;%%
%\cite{Brambilla:1999xf}
%\bibitem{Brambilla:1999xf}
N.~Brambilla {\em et al.\/},
%``Potential NRQCD: An effective theory for heavy quarkonium,''
Nucl.\ Phys.\ B {\bf 566} (2000) 275
[hep-ph/9907240].
%%CITATION = HEP-PH 9907240;%%

%\cite{Luke:1999kz}
\bibitem{Luke:1999kz}
M.~E.~Luke, A.~V.~Manohar and I.~Z.~Rothstein,
%``Renormalization group scaling in nonrelativistic QCD,''
Phys.\ Rev.\ D {\bf 61} (2000) 074025
[hep-ph/9910209].
%%CITATION = HEP-PH 9910209;%%


%\cite{Hoang:1998nz}
\bibitem{Hoang:1998nz}
A.~H.~Hoang {\em et al.\/},
%``Quarkonia and the pole mass,''
Phys.\ Rev.\ D {\bf 59} (1999) 114014
[hep-ph/9804227];
%%CITATION = HEP-PH 9804227;%%
%\bibitem{Brambilla:1999xf}
N.~Brambilla {\em et al.\/},
%``Potential NRQCD: An effective theory for heavy quarkonium,''
Nucl.\ Phys.\ B {\bf 566} (2000) 275
[hep-ph/9907240];
%%CITATION = HEP-PH 9907240;%%
%\bibitem{Beneke:1998rk}
M.~Beneke,
%``A quark mass definition adequate for threshold problems,''
Phys.\ Lett.\ B {\bf 434} (1998) 115
[hep-ph/9804241].
%%CITATION = HEP-PH 9804241;%%

%\cite{Neubert:1994wq}
\bibitem{Neubert:1994wq}
M.~Neubert and C.~T.~Sachrajda,
 %``Cancellation of renormalon ambiguities in the heavy quark effective
%theory,''
Nucl.\ Phys.\ B {\bf 438} (1995) 235
[hep-ph/9407394];
%%CITATION = HEP-PH 9407394;%%
%\bibitem{Luke:1994xd}
M.~E.~Luke, A.~V.~Manohar and M.~J.~Savage,
%``Renormalons in effective field theories,''
Phys.\ Rev.\ D {\bf 51} (1995) 4924
[hep-ph/9407407];
%%CITATION = HEP-PH 9407407;%%
%\bibitem{Grozin:1997ih}
A.~G.~Grozin and M.~Neubert,
%``Higher-order estimates of the chromomagnetic moment of a heavy quark,''
Nucl.\ Phys.\ B {\bf 508} (1997) 311
[hep-ph/9707318].
%%CITATION = HEP-PH 9707318;%%

\bibitem{Brambilla:1999qa}
N.~Brambilla {\em et al.\/},
%``The in_frared behaviour of the static potential 
%in perturbative {QCD},''
Phys.\ Rev.\ D {\bf 60} (1999) 091502
[hep-ph/9903355];
%%CITATION = HEP-PH 9903355;%%

%\cite{Pineda:2001zq}
\bibitem{Pineda:2001zq}
A.~Pineda,
%``Determination of the bottom quark mass from the Upsilon(1S) system,''
JHEP {\bf 0106} (2001) 022
[hep-ph/0105008].
%%CITATION = HEP-PH 0105008;%%

%\cite{Lee:2003hh}
\bibitem{Lee:2003hh}
T.~Lee,
 %``Heavy quark mass determination from the quarkonium ground state energy:  A
%pole mass approach,''
JHEP {\bf 0310} (2003) 044
[hep-ph/0304185].
%%CITATION = HEP-PH 0304185;%%


\bibitem{CCG}
C.~Contreras, G.~Cveti\v c and P.~Gaete,
%``Calculations of binding energies and masses of heavy quarkonia using
%renormalon cancellation,''
Phys.\ Rev.\ D {\bf 70} (2004) 034008
[hep-ph/0311202].
%%CITATION = HEP-PH 0311202;%%

%\cite{Cvetic:2003wk}
\bibitem{Cvetic:2003wk}
G.~Cveti\v c,
%``Estimate of the three-loop contribution to the QCD static potential 
%from renormalon cancellation,''
J.\ Phys.\ G {\bf 30} (2004) 863
[hep-ph/0309262].
%%CITATION = HEP-PH 0309262;%%

%\cite{Caprini:1998wg}
\bibitem{Caprini:1998wg}
I.~Caprini and J.~Fischer,
%``Accelerated convergence of perturbative {QCD} by 
%optimal con_formal  mapping of the Borel plane,''
Phys.\ Rev.\ D {\bf 60} (1999) 054014
[hep-ph/9811367].
%%CITATION = HEP-PH 9811367;%%

%\cite{Cvetic:2001sn}
\bibitem{Cvetic:2001sn}
G.~Cveti\v c and T.~Lee,
%``Bilocal expansion of Borel amplitude 
%and hadronic tau decay width,''
Phys.\ Rev.\ D {\bf 64}, 014030 (2001)
[hep-ph/0101297];
%%CITATION = HEP-PH 0101297;%%
%\cite{Cvetic:2001pg}
%\bibitem{Cvetic:2001pg}
G.~Cveti\v c, T.~Lee and I.~Schmidt,
%``Resummations with renormalon effects for the leading hadronic  
%contribution to the muon (g-2),''
Phys.\ Lett.\ B {\bf 520} (2001) 222
[hep-ph/0107069];
%%CITATION = HEP-PH 0107069;%%

%\cite{Cvetic:2002qf}
\bibitem{Cvetic:2002qf}
G.~Cveti\v c,
%``Borel integration prescription suggested in 
%pQCD by in_frared renormalons,''
Phys.\ Rev.\ D {\bf 67} (2003) 074022
[hep-ph/0211226].
%%CITATION = HEP-PH 0211226;%%

%\cite{Gupta:pd}
\bibitem{Gupta:pd}
S.~N.~Gupta and S.~F.~Radford,
%``Quark Quark And Quark - Anti-Quark Potentials,''
Phys.\ Rev.\ D {\bf 24} (1981) 2309;
%%CITATION = PHRVA,D24,2309;%%
%\cite{Gupta:qc}
%\bibitem{Gupta:qc}
%S.~N.~Gupta and S.~F.~Radford,
%``Remarks On Quark Quark And Quark - Anti-Quark Potentials,''
Phys.\ Rev.\ D {\bf 25} (1982) 3430.
%%CITATION = PHRVA,D25,3430;%%


%\cite{Titard:1993nn}
\bibitem{Titard:1993nn}
S.~Titard and F.~J.~Yndurain,
%``Rigorous QCD evaluation of spectrum and ground state 
%properties of heavy q anti-q systems: With a precision 
%determination of m(b) M(eta(b)),''
Phys.\ Rev.\ D {\bf 49} (1994) 6007
[hep-ph/9310236];
%%CITATION = HEP-PH 9310236;%%
%\cite{Pineda:1997hz}
%\bibitem{Pineda:1997hz}
A.~Pineda and F.~J.~Yndurain,
%``Calculation of quarkonium spectrum and 
%m(b), m(c) to order  alpha(s)**4,''
Phys.\ Rev.\ D {\bf 58} (1998) 094022
[hep-ph/9711287];
%%CITATION = HEP-PH 9711287;%%
%\cite{Pineda:1998ja}
%\bibitem{Pineda:1998ja}
%A.~Pineda and F.~J.~Yndurain,
%``Comment on 'Calculation of quarkonium spectrum 
%and m(b), m(c) to order  alpha(s)**4',''
Phys.\ Rev.\ D {\bf 61} (2000) 077505
[hep-ph/9812371].
%%CITATION = HEP-PH 9812371;%%

%\cite{Czarnecki:1997vz}
\bibitem{Czarnecki:1997vz}
A.~Czarnecki and K.~Melnikov,
%``Two-loop QCD corrections to the heavy quark pair 
%production cross  section in e+ e- 
%annihilation near the threshold,''
Phys.\ Rev.\ Lett.\  {\bf 80} (1998) 2531
[hep-ph/9712222];
%%CITATION = HEP-PH 9712222;%%
%\cite{Czarnecki:2001gi}
%\bibitem{Czarnecki:2001gi}
%A.~Czarnecki and K.~Melnikov,
%``Top quark threshold production at a gamma gamma 
%collider at next-to-next-to-leading order,''
Phys.\ Rev.\ D {\bf 65} (2002) 051501
[hep-ph/0108233];
%%CITATION = HEP-PH 0108233;%%
%\cite{Hoang:1998xf}
%\bibitem{Hoang:1998xf}
A.~H.~Hoang and T.~Teubner,
%``Top quark pair production at threshold: 
%Complete  next-to-next-to-leading order 
%relativistic corrections,''
Phys.\ Rev.\ D {\bf 58} (1998) 114023
[hep-ph/9801397];
%%CITATION = HEP-PH 9801397;%%
%\cite{Kuhn:1998uy}
%\bibitem{Kuhn:1998uy}
J.~H.~K\"uhn, A.~A.~Penin and A.~A.~Pivovarov,
%``Coulomb resummation for b anti-b system near threshold and 
%precision  determination of alpha(s) and m(b),''
Nucl.\ Phys.\ B {\bf 534} (1998) 356
[hep-ph/9801356].
%%CITATION = HEP-PH 9801356;%%

%\cite{Melnikov:1998ug}
\bibitem{Melnikov:1998ug}
K.~Melnikov and A.~Yelkhovsky,
%``The b quark low-scale running mass from Upsilon sum rules,''
Phys.\ Rev.\ D {\bf 59} (1999) 114009
[hep-ph/9805270].
%%CITATION = HEP-PH 9805270;%%

%\cite{Penin:1998zh}
\bibitem{Penin:1998zh}
A.~A.~Penin and A.~A.~Pivovarov,
%``Next-to-next-to-leading order vacuum polarization function of 
%heavy  quark near threshold and sum rules for b anti-b system,''
Phys.\ Lett.\ B {\bf 435} (1998) 413
[hep-ph/9803363];
%%CITATION = HEP-PH 9803363;%%
%\cite{Penin:1998kx}
%\bibitem{Penin:1998kx}
%A.~A.~Penin and A.~A.~Pivovarov,
%``Bottom quark pole mass and $|$V(cb)$|$ matrix element 
%from R(e+ e- $\to$ b  anti-b) and 
%Gamma(sl)(b $\to$ c l nu/l) in the
%next-to-next-to-leading  order,''
Nucl.\ Phys.\ B {\bf 549} (1999) 217
[hep-ph/9807421].
%%CITATION = HEP-PH 9807421;%%

%\cite{Kniehl:2002br}
\bibitem{Kniehl:2002br}
B.~A.~Kniehl {\em et al.\/},
%``Potential NRQCD and heavy-quarkonium spectrum at  
%next-to-next-to-next-to-leading order,''
Nucl.\ Phys.\ B {\bf 635} (2002) 357
[hep-ph/0203166].
%%CITATION = HEP-PH 0203166;%%

%\cite{Penin:2002zv}
\bibitem{Penin:2002zv}
A.~A.~Penin and M.~Steinhauser,
%``Heavy quarkonium spectrum at O(alpha(s)**5 m(q)) 
%and bottom / top quark  mass determination,''
Phys.\ Lett.\ B {\bf 538} (2002) 335
[hep-ph/0204290].
%%CITATION = HEP-PH 0204290;%%

\bibitem{ECH}
G.~Grunberg, Phys. Lett. {\bf 95B} (1980) 70,
{\bf 110B} (1982) 501(E);
%%CITATION = PHLTA,B95,70;%%
{\bf 114B} (1982) 271.
%%CITATION = PHLTA,B114,271;%%


%\cite{Ioffe:2002be}
\bibitem{Ioffe:2002be}
B.~L.~Ioffe and K.~N.~Zyablyuk,
%``Gluon condensate in charmonium sum 
%rules with 3-loop corrections,''
Eur.\ Phys.\ J.\ C {\bf 27} (2003) 229
[hep-ph/0207183].
%%CITATION = HEP-PH 0207183;%%

%\cite{Brambilla:2001qk}
\bibitem{Brambilla:2001qk}
N.~Brambilla, Y.~Sumino and A.~Vairo,
%``Quarkonium spectroscopy and perturbative QCD: 
%Massive quark loop  effects,''
Phys.\ Rev.\ D {\bf 65} (2002) 034001
[hep-ph/0108084].
%%CITATION = HEP-PH 0108084;%%

%\cite{Corcella:2002uu}
\bibitem{Corcella:2002uu}
G.~Corcella and A.~H.~Hoang,
%``Uncertainties in the MS-bar bottom quark mass 
%from relativistic sum  rules,''
Phys.\ Lett.\ B {\bf 554} (2003) 133
[hep-ph/0212297].
%%CITATION = HEP-PH 0212297;%%

%\cite{Eidemuller:2002wk}
\bibitem{Eidemuller:2002wk}
M.~Eidem\"uller,
%``QCD moment sum rules for Coulomb systems: 
%The charm and bottom quark  masses,''
Phys.\ Rev.\ D {\bf 67} (2003) 113002
[hep-ph/0207237].
%%CITATION = HEP-PH 0207237;%%

\bibitem{Narison:1994ag}
S.~Narison,
%``A Fresh look into the heavy quark mass values,''
Phys.\ Lett.\ B {\bf 341} (1994) 73
[hep-ph/9408376]; 
%%CITATION = HEP-PH 9408376;%%
also:
%\bibitem{Narison:1987qh}
S.~Narison,
%``Heavy Quark Mass In The Ms Scheme: Revisited,''
Phys.\ Lett.\ B {\bf 197} (1987) 405.
%%CITATION = PHLTA,B197,405;%%


\end{thebibliography}

\documentclass[fleqn,twoside]{article}
\usepackage[headings]{espcrc2}

\readRCS
$Id: espcrc2.tex,v 1.2 2004/02/24 11:22:11 spepping Exp $
\usepackage{graphicx}
\usepackage[figuresright]{rotating}

\newcommand{\AmS}{{\protect\the\textfont2
  A\kern-.1667em\lower.5ex\hbox{M}\kern-.125emS}}

\title{Bottonium mass -- evaluation using renormalon cancellation}

\author{C. Contreras\address[MCSD]{Physics Department, Universidad T\'ecnica Federico Santa Mar\'{\i}a, Valpara\'{\i}so, Chile},
        G. Cveti\v c\addressmark\thanks{The work supported in 
part by FONDECYT grant No.~1010094 (G.C.), and project 
USM No.~110321 of the UTFSM (C.C. and G.C). Talk presented by G.C., at QCD04, Montpellier, France, July 5-9, 2004.},
        and
        P. Gaete\addressmark[MCSD]}
       
\begin{document}

\begin{abstract}
We present a method of calculating the bottonium mass $M_{\Upsilon}(1S)
= 2 m_b\!+\!E_{b \bar b}$. The binding energy is separated into the soft 
and ultrasoft components $E_{b \bar b}=E_{b \bar b}(s)+E_{b \bar b}(us)$
by requiring the reproduction of the correct residue parameter value
of the renormalon singularity for the renormalon cancellation
in the sum $2 m_b\!+\!E_{b \bar b}(s)$.
The Borel resummation is then performed separately for
$2 m_b$ and $E_{b \bar b}(s)$, using the infrared safe
${\overline m}_b$ mass as input. $E_{b \bar b}(us)$ is estimated.
Comparing the result with the measured value of $M_{\Upsilon}(1S)$,
the extracted value of the quark mass is  
${\overline m}_b(\mu\!=\!{\overline m}_b) = 4.241 \pm 0.068$ GeV
(for the central value $\alpha_s(M_Z)=0.1180$).
This value of ${\overline m}_b$ is close to the earlier values 
obtained from the QCD spectral sum rules, but lower than from pQCD 
evaluations without the renormalon structure for heavy quarkonia.
\vspace{1pc}
\end{abstract}

\maketitle


Heavy quarkonia $q \bar q$ ($q=b, t$) can be investigated
by perturbative methods (pQCD) via effective theories
NPQCD \cite{Caswell:1985ui} and pNRQCD \cite{Pineda:1997bj}
(or: vNRQCD \cite{Luke:1999kz}) because of the scale
hierarchies of the problem:
$m_q > m_q\alpha_s(\mu_s) > m_q \alpha_s^2(\mu_{us}) \stackrel{>}{\sim}
\Lambda_{\rm QCD}$. Here, $m_q$ is the (pole) mass of the
quark, $\mu_s\!\sim\!m_q\alpha_s(\mu_s)$ is the
soft, and $\mu_{us}\!\sim\!m_q \alpha_s^2(\mu_{us})$ the
ultrasoft energy. The quarkonium mass is $M_{q \bar q} = 2 m_q +
E_{q \bar q}$, where the binding energy consists of the soft and
ultrasoft regime contributions: $E_{q \bar q} = E_{q \bar q}(s) +
E_{q \bar q}(us)$. A practical problem which appears in the
course of evaluation of $M_{q \bar q}$ is that the perturbative pole 
mass has an inherent ambiguity $\delta m_q \sim \Lambda_{\rm QCD}$
($\sim\!0.1$ GeV) due to the infrared (IR) renormalon singularity
which appears at the value of the Borel transform variable $b\!=\!1/2$
for $m_q/{\overline m}_q$. Here, ${\overline m}_q$ is the infrared safe
(renormalon-free) ${\overline {\rm MS}}$ mass. However, the static
potential $V_{q \bar q}(r)$ has a related ambiguity
$\delta V_{q \bar q}(r)\!\sim\!\Lambda_{\rm QCD}$ such that
$\delta(2 m_q + V_{q \bar q}) = 0$, i.e., the
renormalon singularity cancels for the combined quantity
$2 m_q\!+\!V_{q \bar q}$ \cite{Hoang:1998nz}
(see also \cite{Neubert:1994wq}). The static potential
is a quantity which does not contain ultrasoft regime
contributions \cite{Brambilla:1999qa}. Therefore, $E_{q \bar q}(s)$
contains the entire $V_{q \bar q}$ and kinetic
energy effects, the latter are renormalon-free. Thus, the $b\!=\!1/2$
renormalon singularity of $V_{q \bar q}$ and $E_{q \bar q}(s)$
are equal and hence the singularity must
cancel also in the combination $2 m_q\!+\!E_{\rm q \bar q}(s)$
\begin{equation}
\delta\left[ 2 m_q + E_{\rm q \bar q}(s) \right] = 0 \ .
\label{rencanc}
\end{equation}  
In principle, $E_{q \bar q}(us)$ could be included in
this relation. However, in practice, it distorts the
cancellation effects since we know these quantities only
to a finite order in perturbation expansions
[cf.~the discussion following Eq.~(\ref{kappa})].

One possibility of evaluating the bottonium ground state
$\Upsilon(1S)$ mass is to use an infrared-safe (renormalon-free)
quark mass (${\overline m}_b$, $m_b^{\rm RS}$, etc.) and
a common couplant $a(\mu) = \alpha_s(\mu)/\pi$ as inputs
in the evaluation of the available truncated perturbation expansion
(TPS) for $M_{\Upsilon}(1S) = 2 m_b + E_{b \bar b}$, in order
to avoid the $b\!=\!1/2$ renormalon (divergence) problems throughout
\cite{Pineda:2001zq}, and then extract the value of ${\overline m}_b$
from the measured value $M_{\Upsilon}(1S) = 9460$ MeV.

Another possibility is to evaluate $E_{b \bar b}$ in terms of the
pole mass $m_b$ and of $a(\mu)$, taking into account the
$b=1/2$ singularity of $E_{b \bar b}$ (using, e.g., the Principal Value
[PV] prescription in the Borel integration), and adding $2 m_b$ to
$E_{b \bar b}$. From $M_{\Upsilon}(1S) = 
2 m_b^{\rm (PV)}\!+\!E_{b \bar b}(m_b^{({\rm PV})})$,
the PV-value of the pole mass $m_b$ is then extracted, and subsequently
the value of ${\overline m}_b$ (via PV Borel integration prescription).
This is the approach of Ref.~\cite{Lee:2003hh}.

Our approach \cite{CCG} follows to a significant degree the 
latter approach, but with some important modifications:
\begin{enumerate}
\item
The input parameter is the renormalon-free mass
${\overline m}_b \equiv {\overline m}_b(\mu\!=\!{\overline m}_b)$ 
(and, of course, the QCD couplant $a(\mu)$).
The pole mass $m_b=m_b({\overline m}_b;\mu_m)$ is evaluated via Borel
integration, accounting for the $b=1/2$ singularity,
and using a hard renormalization scale $\mu_m\!\sim\!m_b$.
The residue parameter $N_m$ of the $b\!=\!1/2$ singularity is
evaluated from the available TPS for $m_b/{\overline m}_b$.
\item
On the basis of the knowledge of $N_m$, we separate
the binding energy into the soft ($s$) and ultrasoft ($us$) regime
contributions: $E_{b \bar b} = 
E_{b \bar b}(s; \mu_f)\!+\!E_{b \bar b}(us; \mu_f)$, 
where the $s$-$us$ factorization
scale $\mu_f$ parametrizes the separation. The separation is
performed by accounting for the renormalon cancellation in the sum 
$2 m_b\!+\!E_{b \bar b}(s)$: $N_m(m_b)\!=\!N_m(E_{b \bar b}(s;\mu_f))$.
The latter relation fixes $\mu_f$ and thus the separation.
\item
The soft binding energy $E_{b \bar b}(s;\mu_f)$ is then
evaluated via Borel integration, accounting for the $b=1/2$
singularity (using the same prescription, e.g. PV, as for $m_b$),
and using a soft renormalization scale $\mu_s\!\sim\!m_b\alpha_s$.
\item
The value of the ultrasoft part $E_{b \bar b}(us; \mu_f)$ 
is estimated.
\item
{}From $2 m_b + E_{b \bar b}(s) + E_{b \bar b}(us) = M_{\Upsilon}(1S)$,
the value of ${\overline m}_b$ is extracted.
\end{enumerate}
For details, we refer to Ref.~\cite{CCG}.

\section{Evaluation of $m_b$ and $N_m$}
\label{sec:mb}

This part has been performed mostly in 
Refs.~\cite{Pineda:2001zq,Lee:2003hh,Cvetic:2003wk}.
The pole mass is known to NLO:
\begin{equation}
S \equiv \frac{m_b}{{\overline m}_b} - 1 = \frac{4}{3}a(\mu_m)
\sum_{j=0}^{\infty} a^j(\mu_m) r_j(\mu_m) \ ,
\label{mbexp}
\end{equation}
where $r_1$ and $r_2$ are known coefficients ($r_0=1$), e.g., in the
${\overline {\rm MS}}$ scheme, and they
depend on $(\mu_m/{\overline m}_b)$; $\mu_m\!\sim\!{\overline m}_b$. 
The Borel transform is
\begin{eqnarray}
\lefteqn{
B_S(b) = \frac{4}{3} \left[ 1 + \frac{r_1}{1! \beta_0} b +
\frac{r_2}{2! \beta_0^2} b^2 + {\cal O}(b^3) \right] }
\label{Bmb1}
\\ 
& = & \frac{ N_m \pi \mu_m}{{\overline m}_b (1\!-\!2 b)^{1\!+\!\nu}}\!
\sum_{k=0}^{\infty} {\tilde c}_k (1\!-\!2 b)^k\!+\!B_S^{\rm (an.)}(b) ,
\label{Bmb2}
\end{eqnarray}
where $\beta_0\!=\!(11\! -\! 2 n_f/3)/4$, 
$\beta_1\! =\! (102\! -\! 38 n_f/3)/16$, 
$\nu\!=\!\beta_1/(2 \beta_0^2)$ ($n_f\!=\!4$); 
${\tilde c}_0\!=\!1$ and the next three coefficients ${\tilde c}_k$
are known (\cite{Pineda:2001zq} for $k\!=\!1,2$; \cite{CCG} for $k\!=\!3$).
$B_S^{\rm (an.)}(b)$ is the analytic part in the bilocal
expansion (\ref{Bmb2}) \cite{Lee:2003hh}, and it is
known up to $\sim\!b^2$. The residue parameter $N_m$ in Eq.~(\ref{Bmb2})
can be obtained with high precision 
\cite{Pineda:2001zq,Lee:2003hh,Cvetic:2003wk}
\begin{equation}
N_m = \frac{{\overline m}_q}{\mu_m} \frac{1}{\pi} R_S(b=1/2) \ ,
\label{Nmform}
\end{equation}
where, according to (\ref{Bmb2})
\begin{equation}
R_S(b; \mu_m) \equiv  (1 - 2 b)^{1 + \nu} B_{S}(b; \mu_m) \ .
\label{RSm}
\end{equation}
Applying the Pad\'e P[1/1] to the known NNLO TPS of
$R_S(b)$ then gives
\begin{equation}
N_m(n_f\!=\!4) = 0.555 \pm 0.020 \ .
\label{Nmnf4}
\end{equation}
The pole mass $m_b$, with ${\overline m}_b$ and $a(\mu_m)$
as input, can now be evaluated by Borel integration using
the bilocal expression (\ref{Bmb2})
\begin{equation}
S(b) = \frac{1}{\beta}_0 {\rm Re}\! 
\int\! db 
\exp \left( - \frac{b}{\beta_0 a(\mu_m)} \right) B_S(b; \mu_m) \ ,
\label{BSint}
\end{equation}
where the integration path can be taken along a ray in the
first or fourth quadrant 
(the generalized PV prescription 
\cite{Caprini:1998wg,Cvetic:2001sn,Cvetic:2002qf}).

\section{Separation} 
\label{sec:sep}

The TPS of the binding energy $E_{b \bar b}$
\begin{equation}
E_{b \bar b} = - \frac{4 \pi^2}{9} {\overline m}_b a^2 \sum_{k=0}^{\infty}
a^k f_k \ ,
\label{Ebbexp}
\end{equation}
is known to the impressive order ${\cal O}(m_b a^5)$
\cite{Gupta:pd,Titard:1993nn,Czarnecki:1997vz,Melnikov:1998ug,Penin:1998zh,Kniehl:2002br,Penin:2002zv},
i.e., in Eq.~(\ref{Ebbexp}) $f_k$ ($k=1,2,3$) are known ($f_0=1$). 
The renormalization scale used in
expansion (\ref{Ebbexp}) should be soft ($\mu_s\!\sim\!m_b\alpha_s$)
or lower. The ultrasoft contributions appear for the first
time at $\sim\!m_b a^5$ \cite{Kniehl:2002br,Penin:2002zv}, 
i.e., $f_3\!=\!f_3(s)\!+\!f_3(us)$. The
$us$ coefficient can be written as \cite{CCG}: 
$f_3(us)/\pi^3 =27.5\!+\!7.1 \ln \alpha_s(\mu_s)\! -\!14.2 \ln \kappa$,
where $\kappa\!\sim\!1$ is the parameter of the $s$-$us$
factorization scale $\mu_f$: $\mu_f\! =\!\kappa m_b \alpha_s(\mu_s)^{3/2}$.
It can be fixed by the requirement of the renormalon
cancellation in $2 m_b\!+\!E_{b \bar b}(s)$:
\begin{equation}
N_m = \frac{2 \pi}{9} \frac{ {\overline m}_b a(\mu_s) }{\mu_s}
R_{F(s)}(b;\mu_s;\mu_f) {\big |}_{b=1/2} \ ,
\label{NmEqq}
\end{equation}
where, in analogy with $R_S$ of (\ref{RSm}) 
\begin{equation}
R_{F(s)}(b; \mu_s; \mu_f) = (1 - 2 b)^{1 + \nu} B_{F(s)}(b;\mu_s; \mu_f) \ ,
\label{RFs}
\end{equation}
and $B_{F(s)}$ is the Borel transform of the quantity
$F(s) = - (9/(4 \pi^2)) E_{b \bar b}(s)/({\overline m}_b a(\mu_s))$
[in analogy with $S$ of (\ref{mbexp})]. Since now the TPS of
$R_{F(s)}$ is known to $\sim\!b^3$, the Pad\'e ${\rm P[2/1]}(b)$
thereof can be taken; using then the value (\ref{Nmnf4})
of $N_m$, the renormalon cancellation condition (\ref{NmEqq})
gives numerically the $s$-$us$ separation parameter
\begin{equation}
\kappa = 0.59 \pm 0.19 \ .
\label{kappa}
\end{equation}
It was possible to obtain the value of $\mu_f$ ($\Leftrightarrow \kappa$)
because the dependence on $\mu_f$ in $N_m$ of Eq.~(\ref{NmEqq}) was taken
(and is known) only to the leading order, the ultrasoft part
was excluded, and $N_m$ is well-known (\ref{Nmnf4}). 
This is similar to the scale-fixing in the
effective charge (ECH) method \cite{ECH}.
If the ultrasoft contributions are included in Eq.~(\ref{NmEqq}),
the value (\ref{Nmnf4}) of $N_m$ cannot be reproduced. 

\section{Evaluation of the soft contributions}
\label{sec:Esoft}

Knowing now the expansion of 
$F(s) = - (9/(4 \pi^2)) E_{b \bar b}(s)/({\overline m}_b a(\mu_s))$
up to $\sim\!a^4$, the Borel transform of this
quantity can be constructed, e.g., with the approach of
the ``$\sigma$-regularized'' bilocal expansion \cite{CCG},
which is a generalization of the bilocal expansion (\ref{Bmb2})
\begin{eqnarray}
\lefteqn{
B_{F(s)}(b) = 
\frac{ 9 N_m  \mu_s}{2 \pi {\overline m}_b a(\mu_s) 
(1\!-\!2 b)^{1\!+\!\nu}}\!
\left[ \sum_{k=0}^{\infty} {\tilde {\cal C}}_k (1\!-\!2 b)^k \right] }
\nonumber\\
&& \times \exp\left[ - \frac{1}{8 \sigma^2} (1\!-\!2 b)^2 \right] +
B_{F(s)}^{\rm (an.)}(b) \ .
\label{Bfs}
\end{eqnarray}
The exponential was introduced in order to suppress the
renormalon part away from $b \approx 1/2$.
The first four coefficients ${\tilde {\cal C}}_k$ are known
(${\tilde {\cal C}}_0=1$), and the analytic part is known now up to
$\sim\!b^3$. The analytic part we can
evaluate either as TPS or as Pad\'e ${\rm P[2/1]}(b)$.
The requirement of the absence of the pole around
$b\!=\!1/2$ in that part, and the independence (weak dependence)
on the renormalization scale $\mu_s$ for the 
Borel-resummed result $E_{b \bar b}(s)$, lead us to
fix the $\sigma$ parameter to the values $\sigma = 0.36 \pm 0.03$.
The Borel integration is performed as in
Eq.~(\ref{BSint}), with the ray (PV) path prescription taken. 

\section{Estimate of the ultrasoft contribution}
\label{sec:Eus}
The ultrasoft part of the energy is known
only to the leading order ($\sim\!m_b a^5$)
\begin{eqnarray}
E_{b \bar b}(us)^{\rm (p)} &\approx& 
- \frac{4}{9} {\overline m}_q \pi^2 f_{3}(us) a^5(\mu_{us})
\nonumber\\
&\approx & (- 150 \pm 100) \ {\rm MeV} \ .
\label{Ebbusp}
\end{eqnarray}
Here, $f_3(us; \mu_f)$ was determined in Sec.~\ref{sec:sep};
the ultrasoft renormalization scale $\mu_{us}$ should be
$\sim\!\alpha_s^2 m_b$, but was taken
numerically to be higher, in the soft regime
($\mu \approx 1.5$-$2.0$ GeV $\ \Rightarrow
\ \alpha_s(\mu) \approx 0.30$-$0.35$), because perturbative
QCD does not allow a running to very low scales.
The bottom mass value was taken ${\overline m}_b = 4.2$ GeV.
The nonperturbative contribution comes primarily from the
gluonic condensate and gives $E_{b \bar b}(us)^{\rm (np)}
\approx 50 \pm 35$ MeV if the gluon condensate values
$\langle (\alpha_s/\pi) G^2 \rangle =
0.009 \pm 0.007 \ {\rm GeV}^4$  \cite{Ioffe:2002be}
are taken. This then results in the following estimate
of the ultrasoft contributions to the binding energy
\begin{equation}
E_{b \bar b}(us)^{\rm (p+np)} \approx (-100 \pm 106) \ {\rm MeV} \ .
\label{Ebbus}
\end{equation} 
In addition, there are contributions to the 
$\Upsilon(1S)$ mass 
due to the nonzero mass of the charm quark
\cite{Brambilla:2001qk}
$\delta M_{\Upsilon}(1S, m_c \not=0) \approx 25 \pm 10$ MeV.

\section{Extraction of the mass ${\overline m}_b$}
\label{sec:concl}

Adding together the Borel-resummed values $2 m_b$,
$E_{b \bar b}(s)$ and $E_{b \bar b}(us)$,
requiring the reproduction of the measured mass
value $M_{\Upsilon}(1S)$ (with the mentioned
$m_c\!\not=\!0$ effect subtracted), 
we extract the following value
for the mass ${\overline m}_b \equiv
{\overline m}_b(\mu\!=\!{\overline m}_b)$:
\begin{equation}
{\overline m}_b({\overline m}_b) = 4.241 \pm 0.068 \ {\rm GeV} \ ,
\label{mbbav2}
\end{equation}
when the QCD coupling value is taken as
$\alpha_s(M_Z) = 0.1180 \pm 0.0015$.
The major source of uncertainty in the result
(\ref{mbbav2}) is the uncertainty from the
ultrasoft contributions (\ref{Ebbus})
($\pm 0.049$ GeV). The other appreciable uncertainties
are from the ambiguity of the soft renormalization
scale $\mu_s\!=\!3\!\pm\!1$ GeV ($\pm 0.013$ GeV) and
of $\alpha_s(M_Z) = 0.1180 \pm 0.0015$
($\pm 0.013$ GeV). If the central value of the gluon
condensate $\langle (\alpha_s/\pi) G^2 \rangle$
is increased from $0.009$ \cite{Ioffe:2002be} to
$0.024 \ {\rm GeV}^4$ 
(used in \cite{Corcella:2002uu,Eidemuller:2002wk}), 
the central value (\ref{mbbav2}) decreases
to ${\overline m}_b({\overline m}_b) = 4.204$ GeV.
This is close to the values of QCD spectral sum rule
calculations which gave central values
${\overline m}_b({\overline m}_b) = 4.20$ GeV
\cite{Corcella:2002uu}; $4.24$ GeV \cite{Eidemuller:2002wk};
and ${\overline m}_b(m_b)=4.23$ GeV \cite{Narison:1994ag}
(Ref.~\cite{Narison:1994ag} uses central condensate value
$\langle (\alpha_s/\pi) G^2 \rangle =0.019$).
The TPS evaluation of $M_{\Upsilon}(1S)$,
without accounting for the renormalon
problem, extracts higher central values 
${\overline m}_b({\overline m}_b) = 4.349$ 
\cite{Penin:2002zv}.

\end{document}